# Temperature Dependence of Heavy Quark Diffusion from (2+1)-flavor Lattice QCD

### v1

@@ 4.3 Chromo-magnetic correlators and spectral functions at NLO
Figure 11: Lattice results for $G_{B}$ normalized by the NLO result for $\mu=4$ GeV.
The chromo-magnetic correlator has also been calculated at NLO [27, 22, 21, 54]. The corresponding spectral function at $\omega\ll T$ has the form:
@@
However, unlike in the chromo-electric case, $\kappa_{B}$ cannot be calculated perturbatively. This is because $\kappa_{B}$ depends on the non-perturbative chromo-magnetic screening mass $m_{M}\sim g^{2}T$ [27].
For $\omega\gg T$, the chromo-magnetic spectral function has the form [27, 22, 21, 54]:
$\rho_{B}^{\rm{UV},{\rm NLO}}(\omega,T,\mu)$ $=$ $\frac{g^{2}_{{\overline{\rm MS}}}(\mu)C_{F}}{6\pi}\omega^{3}\left\{1+\frac{g^{2}_{{\overline{\rm MS}}}(\mu)}{(4\pi)^{2}}\left[-N_{f}\left(\frac{2}{3}\ln\frac{\mu^{2}}{4\omega^{2}}+\frac{26}{9}\right)\right.\right.+$
$\hskip 142.26378pt\left.\left.N_{c}\left(\frac{5}{3}\ln\frac{\mu^{2}}{4\omega^{2}}+\frac{134}{9}-\frac{2\pi^{2}}{3}\right)\right]\right\}\,.$
Unlike $\rho_{E}$, it depends on the renormalization scale $\mu$ already at this order. Note that there was a mistake in the original result [22], which was corrected in [21, 54]. This is a different mistake than the one found in the chromo-electric spectral function, even though both come from the same source. We can reconstruct $G_{B}^{{\rm NLO}}$ from the above spectral function. In Fig. 11, we show the lattice results for $G_{B}$ normalized by $G_{B}^{{\rm NLO}}$ for $\mu=4$ GeV. As one can see from the figure, the NLO result can capture the behavior of $G_{B}$ at small $\tau$, at least for sufficiently high temperatures.
@@ 5.1 Determination of $\kappa_{E}$
Figure 14: Left: Temperature dependence of $\kappa_{E}/T^{3}$ obtained from the combined analysis of different fits, see text. We also show the values of $\kappa_{E}$ obtained in quenched QCD [19, 17]. Right: $K_{E}$ obtained at different temperatures for smax Ansatz.
For small $\tau$, the NLO result with running coupling constant $G_{E}^{\rm NLO+}$ describes the lattice results for the chromo-electric correlator $G_{E}$, c.f. Fig. 8. Therefore, the NLO perturbative result for $\rho_{E}$ is a good approximation of the spectral function. Furthermore, we have seen that combining $\rho_{E}^{\rm IR}(\omega,T)$ and $\rho_{E}^{\rm UV,NLO}(\omega)$ with the diffusion coefficient fixed by with $\kappa_{E}^{\rm NLO}$ in the simple way of (4.14), results in a spectral function that can describe fairly well the chromo-electric correlator at our highest three temperatures, c.f. Fig. 10. Thus, our Ansatz for $\rho_{E}(\omega,T)$ can be some combination of $\rho_{E}^{\rm IR}(\omega)$ and $\rho_{E}^{\rm UV,NLO}(\omega)$ or $\rho_{E}^{\rm LO}(\omega)$. We use the latter to check the sensitivity of $\kappa_{E}$ to the precise form of the perturbative determination of the spectral function at large $\omega$. To fully account for missing higher order perturbative corrections, we use $\rho_{E}^{\rm UV}(\omega)=K_{E}\rho_{E}^{\rm UV,NLO(LO)}(\omega)$, with $K_{E}$ an extra fit parameter. As in previous works [19, 20, 21, 52], we will use the following forms:
@@
The power law Ansatz has two additional parameters, $\omega_{\rm{IR}}$ and $\omega_{\rm{UV}}$. Following [20], and based on known physics considerations, we use $\omega_{\rm{IR}}/T=$0.5, 1.0, 2.0 and $\omega_{\rm{UV}}/T=2\pi$. Furthermore, the parameters $c$ and $p$ in this Ansatz are fixed by requiring continuity of the spectral function at $\omega=\omega_{\rm{IR}}$ and $\omega=\omega_{\rm UV}$. The max Ansatz and the power-law Ansatz do not correspond to smooth spectral functions. This is not a problem, since the fits to the lattice results are only sensitive to the integrals of the spectral function. Furthermore, the non-smoothness of the above Ansätze can be fixed by using smooth theta function. For example, the max Ansatz can be written as:
@@
Altogether, we have 12 different Ansätze to fit the lattice results for $G_{E}$, as illustrated in Fig. 12 for two representative temperatures, $T=195$ MeV and $T=400$ MeV. The corresponding fits work well as discussed in the Appendix B. As an example, Fig. 13 displays the values of $\kappa_{E}$ extracted from the various fitting techniques at the same two representative temperatures. To determine the final estimate of $\kappa_{E}$, we generate $1000$ bootstrap samples for each of the 12 Ansätze and compute the mean and the 68% confidence interval from the combined distribution. The resulting temperature dependence of $\kappa_{E}/T^{3}$ is shown in Fig. 14 (left), where we clearly see a decreasing trend with temperature, as expected. The final values are also listed in Table 5. For comparison, Fig. 14 includes the quenched ($N_{f}=0$) results from Refs. [19, 17], converted to physical units using $r_{0}T_{c}=0.7457$ for the deconfinement phase transition temperature [62] and $r_{0}=0.472$. Remarkably, the (2+1)-flavor QCD results and the quenched results for $\kappa_{E}/T^{3}$ are consistent within errors.
In Fig. 14 (right), we also show the extracted values of $K_{E}$ as a function of temperature for the smax Ansatz. Fits using the LO form of $\rho_{E}^{\rm UV}$ yield $K_{E}$ values that are systematically larger than one by about 20%, while fits using the NLO form gives $K_{E}$ values that approach one at high temperatures. This behavior is expected: at high temperatures, the coupling $g^{2}$ becomes small, making the NLO prediction more accurate and reducing higher-order corrections. In contrast, the LO form lacks sufficient accuracy, even at the smallest $\tau$ probed by our lattice calculations. Although Fig. 14 shows results for $K_{E}$ only for the smax model, the qualitative behavior is similar across all fitting techniques.
@@ 5.2 Determination of $\kappa_{B}$
Figure 16: Left: Temperature dependence of $\kappa_{B}/T^{3}$ obtained from the combined analysis of different fits, see text. Also shown there are the values of $\kappa_{B}$ obtained in quenched QCD [18, 22] Right: $K_{B}$ obtained at different temperatures for smax Ansatz.
@@
Our procedure for determining $\kappa_{B}$ closely follows the one used for $\kappa_{E}$, described above. Specifically, we consider 12 different Ansätze for the spectral function $\rho_{B}(\omega,T)$ that interpolate between the infrared form $\rho_{B}^{\rm IR}(\omega,T)=\kappa_{B}(T)\omega/(2T)$ and the ultraviolet one $\rho_{B}^{\rm UV}=K_{B}\rho_{B}^{\rm LO,NLO}(\omega)$, i.e. use the fit forms given by Eqs. (5.1)-(5.5), but with subscript $E$ being replace by subscript $B$. Both $\kappa_{B}$ and $K_{B}$ are treated as fit parameters. The key difference in this case is that the chromo-magnetic correlator $G_{B}$ depends on the renormalization scale $\mu$ even though $\kappa_{B}$ is independent of the renormalization scale. To control this dependence, we perform the analysis using four values of the renormalization scale: $\mu=2,~{}3,~{}4$ and 5 GeV. These choices are made to avoid large logarithmic corrections in the NLO expression for $\rho_{B}$ of Eq. (4.3). Ideally, the extracted value of $\kappa_{B}$ should be independent of $\mu$. In Fig. 15, we show the results for $\kappa_{B}$ obtained from the different fits at each of the four values of $\mu$. From the figure it is clear that the dependence on $\mu$ is significantly smaller than both the statistical uncertainties and the split arising from the different fit forms. The fits themselves are also of well behaved, as discussed in Appendix B.
@@
To obtain the final estimates for $\kappa_{B}$, we generate $1000$ bootstrap samples for each combination of fit Ansatz and renormalization scale. The resulting distributions are used to extract central values and uncertainties, where we report the median and the 68% confidence interval as the final result. The temperature dependence of $\kappa_{B}/T^{3}$ is shown in Fig. 16 (left). We see that $\kappa_{B}/T^{3}$ is smaller than $\kappa_{E}/T^{3}$, and furthermore it decreases with increasing temperature, as expected. We also compare our results on $\kappa_{B}/T^{3}$ with the quenched results from Refs. [18, 22], using the same procedure describe above for converting $T/T_{c}$ to physical units. As for the electric diffusion coefficient, the quenched and 2+1-flavor results for $\kappa_{B}/T^{3}$ agree.
@@
In the right panel of Fig. 16, we show the values of $K_{B}$ extracted at various temperatures using the smax Ansatz. When the NLO form of $\rho_{B}^{\rm UV}$ is used, the fitted values of $K_{B}$ are consistent with one within errors. This indicates that the NLO expression provides a reliable description of the large-$\omega$ behavior of the chromo-magnetic spectral function. As for $K_{E}$, although we only present results for $K_{B}$ using the smax model, the qualitative behavior is similar across the other fitting forms.
@@ 5.3.1 Temperature dependence of $\kappa_{E,B}$
Figure 17: Comparison of the coupling-strength dependence of the electric (left) and magnetic (right) diffusion coefficients. The coupling is evaluated at a scale $\mu=2\pi T$. We also show the quenched $N_{f}=0$ results extracted from the literature [19, 17, 18, 22] In the electric case (left plot), the black and red dashed lines are the NLO perturbative results for $N_{f}=0$ and $N_{f}=3$. The blue and pink bands are fits to the lattice data for $N_{f}=0$ and $N_{f}=3$, respectively, using Eqs. (5.6) and (5.7)
As noted above, $\kappa_{E}/T^{3}$ and $\kappa_{B}/T^{3}$ decrease with increasing temperature. At sufficiently high temperatures, we expect this dependence to be determined by the scale dependence of the coupling constant $g$. Therefore, we fix the renormalization scale $\mu$ to be $2\pi T$ and recast the temperature dependence of $\kappa_{E,B}/T^{3}$ shown in the previous subsections, but now as a function of $g^{2}(\mu=2\pi T)$. This is shown in Fig. 17 using double logarithmic scale. It is easy to see that the $g$-dependence of $\kappa_{E,B}/T^{3}$ is compatible with $g^{4}$ and is incompatible with $g^{2}$ dependence. In this figure, we also show the quenched $N_{f}=0$ results form Refs. [19, 22, 18, 17]. For the $N_{f}=0$ we use $r_{0}\Lambda_{\overline{\rm MS}}=0.647$ from FLAG 2024 Review [63], which is based on Refs. [64, 65, 66, 67, 68, 69, 70, 71], and $r_{0}=0.472$ fm. The NLO results for $\kappa_{E}$ with $\mu=2\pi T$ are shown as dashed lines in the figure. The NLO result agrees with the our lattice determination of $\kappa_{E}$ in 2+1-flavor QCD while remaining lower than the lattice results for quenched QCD.
Inspired by the perturbative calculation, we parametrize $\kappa_{E,B}/T^{3}$ as:
@@
$\kappa_{B}(T)=\frac{g^{4}C_{F}T^{3}}{18\pi}\left(\left[N_{c}+\frac{N_{f}}{2}\right]\ln\frac{1}{g^{2}}+c_{B}\right)\,.$ (5.7)
These parametrizations describes the lattice results on $\kappa_{E}$ and $\kappa_{B}$ very well, as can be seen in Fig. 17. From the fits using Eq. (5.6), we get $c_{E}^{N_{f}=3}=-0.10(11)$ and $c_{E}^{N_{f}=0}=1.43(61)$, Therefore, we see that the lattice results for $N_{f}=3$ are much closer to the perturbative NLO predictions than the $N_{f}=0$ lattice results. In the same way, from the fits to lattice results on $\kappa_{B}$ using Eq. (5.7), we get $c_{B}=5.8(1.3)$ and $c_{B}=3.20(27)$ for $N_{f}=0$ and $N_{f}=3$, respectively.

### v2

@@ 4.3 Chromo-magnetic correlators and spectral functions at NLO
Figure 11: Lattice results for $G_{B}$ normalized by the NLO result for $\mu=4$ GeV. For $T=10$ GeV, we show the results for $\mu=50$ GeV.
The chromo-magnetic correlator has also been calculated at NLO Bouttefeux:2020ycy ; Banerjee:2022uge ; Altenkort:2023eav ; delaCruz:2024cix . The corresponding spectral function at $\omega\ll T$ has the form:
@@
However, unlike in the chromo-electric case, $\kappa_{B}$ cannot be calculated perturbatively. This is because $\kappa_{B}$ depends on the non-perturbative chromo-magnetic screening mass $m_{M}\sim g^{2}T$ Bouttefeux:2020ycy . This is the same Linde problem Linde:1980ts ; Gross:1980br that appears in the pressure at 4 loops and in the Debye mass calculations at 2 loops (NLO) Rebhan:1994mx ; Braaten:1994pk . Here it appears already at leading order because the correlation function involves soft chromo-magnetic fields at leading order, while for the pressure and the Debye mass soft chromo-magnetic fields first contribute at higher loop orders. Thus, perturbation theory only gives the leading-log result for $\kappa_{B}$. This has been carefully discussed in Ref. Bouttefeux:2020ycy .
For $\omega\gg T$, the chromo-magnetic spectral function has the form Bouttefeux:2020ycy ; Banerjee:2022uge ; Altenkort:2023eav ; delaCruz:2024cix :
[…]
$\left.\left.N_{c}\left(\frac{5}{3}\ln\frac{\mu^{2}}{4\omega^{2}}+\frac{134}{9}-\frac{2\pi^{2}}{3}\right)\right]\right\}\,.$
Unlike $\rho_{E}$, it depends on the renormalization scale $\mu$ already at this order. Note that there was a mistake in the original result Banerjee:2022uge , which was corrected in Altenkort:2023eav ; delaCruz:2024cix . This is a different mistake than the one found in the chromo-electric spectral function, even though both come from the same source. We can reconstruct $G_{B}^{{\rm NLO}}$ from the above spectral function. In Fig. 11, we show the lattice results for $G_{B}$ normalized by $G_{B}^{{\rm NLO}}$ for $\mu=4$ GeV. As one can see from the figure, the NLO result can capture the behavior of $G_{B}$ at small $\tau$, at least for sufficiently high temperatures.
@@ 5.1 Determination of $\kappa_{E}$
Figure 14: Left: Temperature dependence of $\kappa_{E}/T^{3}$ obtained from the combined analysis of different fits, see text. We also show the values of $\kappa_{E}$ obtained in quenched QCD Brambilla:2020siz ; Banerjee:2022gen . Right: rescaling parameter $K_{E}$ obtained at different temperatures for the smax Ansatz.
For small $\tau$, the NLO result with running coupling constant $G_{E}^{\rm NLO+}$ describes the lattice results for the chromo-electric correlator $G_{E}$, c.f. Fig. 8. Therefore, the NLO perturbative result for $\rho_{E}$ is a good approximation of the spectral function at larger $\omega$. Furthermore, we have seen that combining $\rho_{E}^{\rm UV,NLO}(\omega)$ and $\rho_{E}^{\rm IR}(\omega,T)$, with the diffusion coefficient fixed to $\kappa_{E}^{\rm NLO}$ as described in (4.14), results in a spectral function that can describe fairly well the chromo-electric correlator at our highest three temperatures, c.f. Fig. 10. Thus, our Ansatz for $\rho_{E}(\omega,T)$ can be some combination of $\rho_{E}^{\rm IR}(\omega)$ and $\rho_{E}^{\rm UV,NLO}(\omega)$ or $\rho_{E}^{\rm LO}(\omega)$. We use the latter to check the sensitivity of $\kappa_{E}$ to the precise form of the perturbative determination of the spectral function at large $\omega$. To fully account for missing higher order perturbative corrections, we use $\rho_{E}^{\rm UV}(\omega)=K_{E}\rho_{E}^{\rm UV,NLO(LO)}(\omega)$, with $K_{E}$ a rescaling factor treated as an extra fit parameter. As in previous works Brambilla:2020siz ; Altenkort:2023oms ; Altenkort:2023eav ; Altenkort:2024spl , we will use the following forms:
@@
The power law Ansatz has two additional parameters, $\omega_{\rm{IR}}$ and $\omega_{\rm{UV}}$. Following Altenkort:2023oms , and based on known physics considerations, we use $\omega_{\rm{IR}}/T=$0.5, 1.0, 2.0 and $\omega_{\rm{UV}}/T=2\pi$. Furthermore, the parameters $c$ and $p$ in this Ansatz are fixed by requiring continuity of the spectral function at $\omega=\omega_{\rm{IR}}$ and $\omega=\omega_{\rm UV}$. The max Ansatz and the power-law Ansatz do not correspond to smooth spectral functions. This is not a problem, since the fits to the lattice results are only sensitive to the integrals of the spectral function. Furthermore, the non-smoothness of the above Ansätze can be fixed by using a smoothed theta function. For example, the max Ansatz can be written as:
@@
Altogether, we have 12 different Ansätze to fit the lattice results for $G_{E}$, as illustrated in Fig. 12 for two representative temperatures, $T=195$ MeV and $T=400$ MeV. The corresponding fits work well as discussed in the Appendix B. As an example, Fig. 13 displays the values of $\kappa_{E}$ extracted from the various fit forms at the same two representative temperatures. To determine the final estimate of $\kappa_{E}$, we generate $1000$ bootstrap samples for each of the 12 Ansätze and compute the mean and the 68% confidence interval from the combined distribution. The resulting temperature dependence of $\kappa_{E}/T^{3}$ is shown in Fig. 14 (left), where we clearly see a decreasing trend with temperature, as expected. The final values are also listed in Table 5. For comparison, Fig. 14 includes the quenched ($N_{f}=0$) results from Refs. Brambilla:2020siz ; Banerjee:2022gen , converted to physical units using $r_{0}T_{c}=0.7457$ for the deconfinement phase transition temperature Francis:2015lha and $r_{0}=0.472$ fm. Remarkably, the (2+1)-flavor QCD results and the quenched results for $\kappa_{E}/T^{3}$ are consistent within errors.
In Fig. 14 (right), we also show the extracted values of the rescaling factor $K_{E}$ as a function of temperature for the smax Ansatz. Fits using the LO form of $\rho_{E}^{\rm UV}$ yield $K_{E}$ values that are systematically larger than one by about 20%, while fits using the NLO form gives $K_{E}$ values that approach one at high temperatures. This behavior is expected: at high temperatures, the coupling $g^{2}$ becomes small, making the NLO prediction more accurate and reducing higher-order corrections. In contrast, the LO form lacks sufficient accuracy, even at the smallest $\tau$ probed by our lattice calculations. Although Fig. 14 shows results for $K_{E}$ only for the smax model, the qualitative behavior is similar across all fitting techniques.
@@ 5.2 Determination of $\kappa_{B}$
Figure 16: Left: Temperature dependence of $\kappa_{B}/T^{3}$ obtained from the combined analysis of different fits, see text. Also shown there are the values of $\kappa_{B}$ obtained in quenched QCD Brambilla:2022xbd ; Banerjee:2022uge Right: $K_{B}$ obtained at different temperatures for smax Ansatz. We do not display the results for $T=163$ MeV due to its large errors.
Our procedure for determining $\kappa_{B}$ closely follows the one used for the determination of $\kappa_{E}$ as described above. Specifically, we consider 12 different Ansätze for the spectral function $\rho_{B}(\omega,T)$ that interpolate between the infrared form $\rho_{B}^{\rm IR}(\omega,T)=\kappa_{B}(T)\omega/(2T)$ and the ultraviolet one $\rho_{B}^{\rm UV}=K_{B}\rho_{B}^{\rm UV,NLO(LO)}(\omega)$, i.e., use the fit forms given by Eqs. (5.1)-(5.5), but with the subscript $E$ being replaced by the subscript $B$. Both $\kappa_{B}$ and $K_{B}$ are treated as fit parameters.
As we stated in section 3.4.1, we do not use the physical scheme for $G_{B}$ introduced in Laine:2021uzs ; Bouttefeux:2020ycy and choose to leave the entire calculation in the $\overline{\rm MS}$} scheme. However, an important step remains to be addressed. The chromo-magnetic correlator $G_{B}(\tau)$ in the $\overline{\rm MS}$} scheme depends explicitly on the renormalization scale $\mu$, and as a result, the corresponding spectral function $\rho_{B}(\omega)$ inherits this scheme and scale dependence across all frequencies $\omega$. This dependence arises because the insertion of a magnetic $B_{i}$ field on a temporal Wilson line induces a non-trivial anomalous dimension, making any correlation function constructed from such operators a scheme and scale dependent quantity.
The physically meaningful quantity is not the $\overline{\rm MS}$} renormalized $\kappa_{B}^{{\overline{\rm MS}}}$, but rather its renormalization-group invariant counterpart $\kappa_{B}^{\rm{phys}}$, which is defined through the zero-frequency limit of a heavy-quark current-current correlator. The relation between both the physical and $\overline{\rm MS}$} correlators was worked out in Laine:2021uzs ; Bouttefeux:2020ycy and takes the form of the following multiplicative factor:
$G^{\rm{phys}}_{B}(\tau T)=G^{{\overline{\rm MS}}}(\tau T,\mu)\left[1+g^{2}\gamma_{0}\left(\ln\frac{\mu^{2}}{(4\pi T)^{2}}-2+2\gamma_{\rm E}\right)\right]\,.$ (5.6)
In this work, we apply this correction as a final multiplicative rescaling to the $\kappa_{B}^{{\overline{\rm MS}}}$ values extracted from the $\overline{\rm MS}$} renormalized correlator, yielding the final result for $\kappa_{B}^{\rm{phys}}$, which we will simply denote as $\kappa_{B}$ in the following for brevity. This step removes the residual scale dependence, and restores the renormalization-group invariance up to higher-order corrections. We emphasize that this correction only applies to $\kappa_{B}$, as the chromo-electric correlator does not carry a non-trivial anomalous dimension in this setup.
@@
To control the scale dependence, we perform the analysis using four values of the renormalization scale: $\mu=2,~3,~4$ and 5 GeV, except for $T=10$ GeV, where we use a higher scale of $\mu=50$ GeV. These choices are made to avoid large logarithmic corrections in the NLO expression for $\rho_{B}$ of Eq. (4.3). Ideally, the extracted value of $\kappa_{B}$ should be independent of $\mu$. In Fig. 15, we show the results for $\kappa_{B}$ obtained from the different fits at each of the four values of $\mu$. From the figure it is clear that the dependence on $\mu$ is significantly smaller than both the statistical uncertainties and the split arising from the different fit forms. The fits themselves are also well behaved, as discussed in Appendix B.
@@
To obtain the final estimates for $\kappa_{B}$, we generate $1000$ bootstrap samples for each combination of fit Ansatz and renormalization scale. The resulting distributions are used to extract central values and uncertainties, where we report the median and the 68% confidence interval as the final result. The temperature dependence of $\kappa_{B}/T^{3}$ is shown in Fig. 16 (left). We see that $\kappa_{B}/T^{3}$ is smaller than $\kappa_{E}/T^{3}$, and furthermore it decreases with increasing temperature, as expected. We also compare our results on $\kappa_{B}/T^{3}$ with the quenched results from Refs. Brambilla:2022xbd ; Banerjee:2022uge , using the same procedure described above for converting $T/T_{c}$ to physical units. As for the electric diffusion coefficient, the quenched and 2+1-flavor results for $\kappa_{B}/T^{3}$ agree.
@@
In the right panel of Fig. 16, we show the values of $K_{B}$ extracted at various temperatures using the smax Ansatz. When the NLO form of $\rho_{B}^{\rm UV}$ is used, the fitted values of $K_{B}$ are consistent with one within errors. This indicates that the NLO expression provides a reliable description of the large-$\omega$ behavior of the chromo-magnetic spectral function. Although we only present results for $K_{B}$ using the smax model, the qualitative behavior is similar across the other fitting forms.
@@ 5.3.1 Temperature dependence of $\kappa_{E,B}$
Figure 17: Comparison of the coupling-strength dependence of the electric (left) and magnetic (right) diffusion coefficients. The coupling is evaluated at a scale $\mu=2\pi T$. We also show the quenched $N_{f}=0$ results from the literature Brambilla:2020siz ; Banerjee:2022gen ; Brambilla:2022xbd ; Banerjee:2022uge . In the electric case (left plot), the black and red dashed lines are the NLO perturbative results for $N_{f}=0$ and $N_{f}=3$. The blue and pink bands are fits to the lattice data for $N_{f}=0$ and $N_{f}=3$, respectively, using Eqs. (5.7) and (5.8)
As noted above, $\kappa_{E}/T^{3}$ and $\kappa_{B}/T^{3}$ decrease with increasing temperature. At sufficiently high temperatures, we expect this dependence to be determined by the scale dependence of the coupling constant $g$. Therefore, we fix the renormalization scale $\mu$ to be $2\pi T$ and recast the temperature dependence of $\kappa_{E,B}/T^{3}$ shown in the previous subsections as the dependence on $g^{2}(\mu=2\pi T)$. This is shown in Fig. 17 using double logarithmic scale. It is easy to see that the $g$-dependence of $\kappa_{E,B}/T^{3}$ is compatible with $g^{4}$ and is incompatible with $g^{2}$ dependence. In this figure, we also show the quenched $N_{f}=0$ results from Refs. Brambilla:2020siz ; Banerjee:2022uge ; Brambilla:2022xbd ; Banerjee:2022gen . For the $N_{f}=0$ we use $r_{0}\Lambda_{\overline{\rm MS}}=0.647$ from FLAG 2024 Review FlavourLatticeAveragingGroupFLAG:2024oxs , which is based on Refs. Capitani:1998mq ; Gockeler:2005rv ; Brambilla:2010pp ; Kitazawa:2016dsl ; Ishikawa:2017xam ; DallaBrida:2019wur ; Bribian:2021cmg ; Brambilla:2023fsi , and $r_{0}=0.472$ fm. The NLO results for $\kappa_{E}$ with $\mu=2\pi T$ are shown as dashed lines in the figure. The NLO result agrees with our lattice determination of $\kappa_{E}$ in 2+1-flavor QCD while remaining lower than the lattice results for quenched QCD.
Inspired by the perturbation theory and weak coupling calculations, we parametrize $\kappa_{E,B}/T^{3}$ as:
@@
$\kappa_{B}(T)=\frac{g^{4}C_{F}T^{3}}{18\pi}\left(\left[N_{c}+\frac{N_{f}}{2}\right]\ln\frac{1}{g^{2}}+c_{B}\right)\,$, (5.8)
using data corresponding to the large temperature region $T\leq 352$ MeV. These parametrizations describe the lattice results on $\kappa_{E}$ and $\kappa_{B}$ very well, as can be seen in Fig. 17. From the fits using Eq. (5.7), we get $c_{E}^{N_{f}=3}=-0.12(19)$ and $c_{E}^{N_{f}=0}=1.43(61)$, Therefore, we see that the lattice results for $N_{f}=3$ are much closer to the perturbative NLO predictions than the $N_{f}=0$ lattice results. In the same way, from the fits to lattice results on $\kappa_{B}$ using Eq. (5.8), we get $c_{B}=5.8(1.3)$ and $c_{B}=1.37(22)$ for $N_{f}=0$ and $N_{f}=3$, respectively.